### v4

# MAGPIE: Multilevel–Adaptive–Guided solver for ptychographic phase retrieval

Abstract
We introduce MAGPIE (Multilevel––Adaptive––Guided Ptychographic Iterative Engine), a stochastic multigrid solver for the ptychographic phase-retrieval problem. The ptychographic phase-retrieval problem is inherently nonconvex and ill-posed. To address these challenges, we reformulate the original nonlinear and nonconvex inverse problem as the iterative minimization of a quadratic surrogate model that majorizes the original objective. This surrogate not only ensures favorable convergence properties but also generalizes the Ptychographic Iterative Engine (PIE) family of algorithms. By solving the surrogate model using a multigrid method, MAGPIE achieves substantial gains in convergence speed and reconstruction quality over traditional approaches.
Keywords phase retrieval  $\cdot$ multigrid optimization  $\cdot$ inverse problems  $\cdot$ ptychography  $\cdot$ coherent diffraction imaging
Ptychography is a state-of-the-art coherent diffractive imaging technique that has attracted significant attention for its ability to reconstruct high-resolution, complex-valued images from a series of diffraction patterns. The technique finds applications in materials science [1, 2], biological imaging [3, 4], integrated circuit imaging [5, 6], and X-ray crystallography [7] (see [8] for a comprehensive overview). In a typical ptychographic experiment, a coherent beam is scanned over overlapping regions of an object while recording the resulting diffraction patterns. Since only the intensity of the diffracted waves is measured—and phase information is lost—the reconstruction task becomes a phase retrieval problem, requiring the recovery of both the amplitude and the phase of the object.
Mathematically, ptychographic phase retrieval is formulated as an inverse problem that is inherently nonconvex and ill-posed. To address these challenges, numerous algorithms have been proposed, including the well-known Ptychographical Iterative Engine (PIE) family [9, 10, 11]. This family of algorithms casts the problem in an iterative fashion and stabilizes the reconstruction through adding a quadratic regularization term.
Another feature of the ptychographic phase retrieval problem is its inherent multiscale structure. Coarse and fine features coexist in the to-be-reconstructed object, requiring different levels of numerical accuracy. This naturally calls for a multigrid solver. Multigrid methods have a long history in the numerical solution of partial differential equations, exploiting the inherent multi-scale structure of the underlying problem by solving it on a hierarchy of discretizations [12]. Coarse-grid approximations capture the global structure of the solution, while fine-grid corrections resolve detailed features. In the context of optimization, a multilevel [^1] Throughout this work, the terms “grid” and “level” are used interchangeably. optimization algorithm (termed MG/OPT) [13] has demonstrated significant improvements in convergence performance, while maintaining flexibility over a wide range of applications [14, 15, 16]. Specific to inverse problems, this hierarchical structure also helps mitigate the ill-posedness inherent in many inverse problems [17, 18, 19]. Recent work has adapted MG/OPT to the ptychographic context [20], resulting in reduced computational costs and enhanced convergence performance.
In this paper, we propose MAGPIE (Multilevel–Adaptive–Guided Ptychographic Iterative Engine), a solver that integrates the advantages of both PIE and MG/OPT. As done in PIE, we recast the problem into an iterative minimization problem with each step solving a quadratic surrogate that majorizes the original objective. The majorization effect brings us a favorable convergence rate. Meanwhile, we also deploy a multigrid-based optimization strategy and reconstruct the coarse and fine features of the object level by level. Unlike other multigrid-based optimization methods whose coarse-level models are generically incompatible with those on the fine level, our proposed surrogate naturally builds compatibility across scales, thereby eliminating the correction term typically required by MG/OPT. This compatibility also ensures automatic hyperparameter selection: once parameters are tuned on the finest grid, those for coarser levels are derived automatically, simplifying tuning in the implementation and improving usability. The overall method is run in a stochastic fashion, and the final algorithm delivers nice reconstruction with substantially reduced cost of full-gradient evaluations and accelerated computational performance.
The remainder of this paper is organized as follows. Section 1 presents the ptychographic phase retrieval problem and introduces two main techniques to be deployed: the surrogate modeling deployed in the PIE family algorithm and multigrid optimization strategy; Section 2 introduces our surrogate model, a model that will be shown to majorize the original objective and thus brings in favorable convergence properties; in Section 3 we finally present our ultimate solver: MAGPIE, through designing a multigrid method for the surrogate model. We also examine the iterative performance of the algorithm. In Section 4 we present numerical results.
1 Preliminaries
1.1 2D ptychographic phase retrieval
In 2D ptychographic phase retrieval, the goal is to reconstruct the object of interest $\bm{z}^{\ast}\in\mathbb{C}^{n^{2}}$ (in vectorized form) from its observed diffraction patterns. The data $\bm{d}_{k}\in\mathbb{R}^{m^{2}}$ are collected as phaseless intensity measurements from overlapping scanning regions. Specifically,
$\bm{d}_{k}=|\mathcal{F}(\bm{Q}\odot P_{k}\bm{z}^{\ast})|^{2}+\epsilon_{k},\quad k=1,\ldots,N\,,$
where $N$ is the total number of scans, $P_{k}\in\mathbb{R}^{m^{2}\times n^{2}}$ projects the image onto the $k$-th scanning region, $\bm{Q}\in\mathbb{C}^{m^{2}}$ represents the probe, $\mathcal{F}\in\mathbb{C}^{m^{2}\times m^{2}}$ denotes the two-dimensional discrete Fourier transform operator, $\epsilon_{k}\in\mathbb{R}^{m^{2}}$ corresponds to the noise associated with the $k$-th measurement, $\odot$ denotes the Hadamard (element-wise) product, and $|\cdot|^{2}$ represents the element-wise intensity. The diagonal elements of $P_{k}$ are $1$ for the columns corresponding to the illuminated pixels in the $k$-th scanning region. We also refer to $\bm{Q}\odot P_{k}\bm{z}^{\ast}$ as the $k$-th exit wave. The overlap ratio refers to the fraction of the probe area that overlaps between adjacent scan positions. For example, a 50% overlap ratio means that the probe is shifted by half of its size in each direction when moving between neighboring positions. Higher overlap ratios increase data redundancy and generally improve the stability and convergence of reconstruction algorithms, at the cost of increased measurement acquisition [8]. Figure 1 illustrates the experimental setup and data acquisition process for ptychography.
Figure 1: Experimental setup and data acquisition for ptychography.
The inverse problem consists in inferring the object of interest, $\bm{z}^{*}$, from the collected data $\{\bm{d}_{k}\}_{k=1}^{N}$. This problem is commonly formulated as the following optimization problem:
$\min_{\bm{z}}\Phi(\bm{z})=\frac{1}{2}\sum_{k=1}^{N}\left\||\mathcal{F}(\bm{Q}\odot P_{k}\bm{z})|-\sqrt{\bm{d}_{k}}\right\|_{2}^{2}\,,$
where $\sqrt{\cdot}$ denotes the element-wise square root. Following the terminology in [11] and invoking Parseval’s theorem, the problem can equivalently be written as:
$\min_{\bm{z}}\Phi(\bm{z})$ $=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot P_{k}\bm{z}-\mathcal{R}_{k}(P_{k}\bm{z})\right\|_{2}^{2}\,,$ (1)
$=\sum_{k=1}^{N}\Phi_{k}(\bm{z}_{k})=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right\|_{2}^{2}\,,$
where $\bm{z}_{k}$ and the operator $\mathcal{R}_{k}$ are defined below, and the symbol $\odot$ is omitted for notational convenience.
Definition 1.1 (Revised Exit Wave).
For the $k$-th scanning region, define
$\bm{z}_{k}=P_{k}\bm{z}\,,$
and define the revised exit wave as
$\mathcal{R}_{k}(\bm{z}_{k})=\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta\left(\mathcal{F}(\bm{Q}\odot\bm{z}_{k})\right)\right)\right)\,,$
where $\theta(\cdot)$ denotes the element-wise principal argument [^2] Since $\theta(0)$ is undefined, at index $r$, we set $\left[\theta\left(\mathcal{F}(\bm{Q}\odot\bm{z}_{k})\right)\right]_{r}=0$ when $\left[\mathcal{F}(\bm{Q}\odot\bm{z}_{k})\right]_{r}=0$..
Since the objective function is real-valued but defined on a complex variable, it is not complex-differentiable in the classical sense [21]. A common remedy is to employ the $\mathbb{C}\mathbb{R}$-calculus [22]. To that end, we represent vectors in the complex field, $\bm{z}_{k}\in\mathbb{C}^{m^{2}}$, in the Euclidean space $\mathbb{R}^{2m^{2}}$ as follows:
$\left\{[\bm{x}_{k}^{\top},\bm{y}_{k}^{\top}]^{\top}:\bm{x}_{k}+i\bm{y}_{k}=\bm{z}_{k}\right\}\,,$
so that the objective function can be interpreted as a mapping from $\mathbb{R}^{2m^{2}}$ to $\mathbb{R}$. This representation enables us to compute the complex gradient of $\Phi_{k}$:
$\nabla_{\bm{z}_{k}}\Phi_{k}$ $=\nabla_{\bm{x}_{k}}\Phi_{k}+i\nabla_{\bm{y}_{k}}\Phi_{k}\,,$ (2)
$=\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right)\,,$
(See Appendix A.1 for the derivation). Consequently, the full gradient of  Eqn. (1) is
$\nabla_{\bm{z}}\Phi(\bm{z})$ $=\sum_{k=1}^{N}P_{k}^{\top}\nabla_{\bm{z}_{k}}\Phi_{k}(P_{k}\bm{z})$
$=\sum_{k=1}^{N}P_{k}^{\top}\overline{\bm{Q}}\left(\bm{Q}P_{k}\bm{z}-\mathcal{R}_{k}(P_{k}\bm{z})\right)\,.$
PIE family of algorithms
The PIE family of algorithms [9, 10, 11] is widely employed to minimize the misfit  Eqn. (1). At each iteration, the scanning regions are randomly shuffled and then updated one by one using a regularized quadratic surrogate model until all regions have been updated. While one scanning region is selected and updated, all entries outside that region remain unchanged. Let $\bm{z}_{k}^{(j)}$ denote the estimate for the $k$-th region at the $j$-th iteration [^3] In each iteration, we shuffle the indices of the scanning regions $\{k\}_{k=1}^{N}$ and update them sequentially.. The local update is then performed by
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\bm{z}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)\,.$ (3)
Specifically, we define its corresponding regularized quadratic surrogate function $\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ as
$\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,,$
where $\bm{u}\in\mathbb{R}^{m^{2}}$ is an a priori chosen regularization coefficient and is applied element-wise to $\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}$. Then, $\bm{z}^{(j+1)}$ is obtained as the final output of  Eqn. (3) after sweeping through all scanning regions. Notably, for a given $\bm{z}_{k}^{(j)}$, this objective function is quadratic in $\bm{z}_{k}$.
Correspondingly, the complex gradient of $\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ with respect to $\bm{z}_{k}$ is:
$\nabla_{\bm{z}_{k}}\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\,,$ (4)
(See Appendix A.2 for the derivation). Therefore, we set the gradient to zero and solve for the explicit solution $\bm{z}_{k}^{+}$:
$\bm{u}\left(\bm{z}_{k}^{+}-\bm{z}_{k}^{(j)}\right)$ $=\overline{\bm{Q}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{+}\right)\,,$
$=\overline{\bm{Q}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\left(\bm{z}_{k}^{+}-\bm{z}_{k}^{(j)}+\bm{z}_{k}^{(j)}\right)\right)\,,$
implying ${\text{implying that}}\quad\left(\bm{u}+|\bm{Q}|^{2}\right)\left(\bm{z}_{k}^{+}-\bm{z}_{k}^{(j)}\right)$ $=\overline{\bm{Q}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)\,,$
recovering the local update rule in Eqn. (3).
The choice of $\bm{u}$ significantly influences the reconstruction accuracy. Within the PIE family of algorithms, experiments in [11] reported that the best performance is achieved by setting:
$\bm{u}=\bm{u}^{\texttt{rPIE}}=\alpha\left(\|\bm{Q}\|^{2}_{\infty}-|\bm{Q}|^{2}\right)\,,$
where $\alpha>0$, and $\|\bm{Q}\|^{2}_{\infty}$ is the maximum intensity among the probe’s entries [^4] In this work, $\bm{Q}$ is flattened into a vector in $\mathbb{C}^{m^{2}}$, so its infinity norm $\|\bm{Q}\|_{\infty}$ equals the maximum magnitude among its entries.. This algorithm is termed regularized Ptychographical Iterative Engine (rPIE). The rationale behind this choice is that in regions of the object with low probe illumination ($|\bm{Q}|_{r}\ll 1$), the revised exit wave is more susceptible to noise, necessitating a larger penalty to stabilize the update. Conversely, where the probe strongly illuminates the object, a smaller penalty is applied, reflecting our increased confidence in those measurements. This adaptive weighting not only enhances the robustness of rPIE but also effectively mitigates noise-induced instability. The pseudocode for rPIE is presented in Algorithm 1. However, despite these advantages, rPIE currently lacks a rigorous theoretical foundation in terms of convergence. We reexamine the algorithm from a surrogate-minimization perspective in Section 2.
Algorithm 1 $\texttt{rPIE}(\bm{z}^{(0)},\bm{Q},\{\bm{d}_{k}\}_{k=1}^{N},\alpha)$
0:  Object: $\bm{z}^{(0)}$, probe: $\bm{Q}$, measurements: $\{\bm{d}_{k}\}_{k=1}^{N}$, regularization constant: $\alpha$.
1:  $\bm{u}\leftarrow\alpha\left(\|\bm{Q}\|^{2}_{\infty}-|\bm{Q}|^{2}\right)$
2:  while $\bm{z}^{(j)}$ not converged do
3:     for $k$ randomly drawn from $1$ to $N$ without replacement do
4:        $\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)$
5:     end for
6:  end while
7:  return  $\bm{z}^{(j)}$
1.2 Multilevel optimization algorithm (MG/OPT)
Ptychographic phase retrieval contains an intrinsic resolution hierarchy, making it an ideal candidate for multigrid acceleration. MG/OPT is a multigrid solver for optimization. Given an objective function $f_{h}$ defined on a fine grid, our goal is to solve $\min_{\bm{z}}f_{h}(\bm{z})$. MG/OPT accelerates this process by, when appropriate, lifting the problem to a coarser level where a simpler optimization problem can be solved more efficiently. By communicating information between fine and coarse grids and offloading expensive fine-grid computations onto the cheaper coarse level, the method speeds up overall convergence. It has already demonstrated faster convergence for the ptychographic phase-retrieval problem [20]. Here, we present the two-level MG/OPT algorithm, which employs one fine grid and one coarse grid; the multilevel extension follows naturally. We introduce the following notation:
  • $h$ and $H$ denote the fine grid and coarse grid, respectively. Let $n_{h}$ be the number of fine-grid points per dimension. For simplicity, we assume $n_{h}=2n_{H}$.
  • $f_{H}$ is a manually defined objective function on the coarse grid, intended to capture key features of $f_{h}$.
  • $\bm{I}_{h}^{H}$: Denotes the restriction operator (binning) that downsamples variables from the fine grid to the coarse grid via average pooling. Specifically,
[…]
  • $\bm{I}_{H}^{h}$: Denotes the prolongation operator that interpolates the search direction from the coarse grid to the fine grid. Specifically,
    $\bm{I}_{H}^{h}=4(\bm{I}_{h}^{H})^{\top}\,.$ (6)
We summarize the process of MG/OPT in Algorithm 2.
Algorithm 2 MG/OPT
0:  Given an initial estimate of the solution $\bm{z}_{h}^{0}$ and integers $k_{1},k_{2}\geq 0$ satisfying $k_{1}+k_{2}>0$.
1:  for $j=0,1,\dots$ until convergence do
2:     Pre-smoothing: Apply $k_{1}$ iterations of a convergent optimization algorithm to $\min_{\bm{z}}f_{h}(\bm{z})$ to obtain $\bar{\bm{z}}_{h}$ (with $\bm{z}_{h}^{(j)}$ as the initial guess).
3:     Recursion:
4:        Compute $\bar{\bm{z}}_{H}=\bm{I}_{h}^{H}\,\bar{\bm{z}}_{h}$ and $\bar{\bm{v}}_{H}=\nabla f_{H}(\bar{\bm{z}}_{H})-\bm{I}_{h}^{H}\nabla f_{h}(\bar{\bm{z}}_{h})$.
5:        Minimize (perhaps approximately) the surrogate model
$f_{s}(\bm{z}_{H})=f_{H}(\bm{z}_{H})-\bar{\bm{v}}_{H}^{T}\bm{z}_{H}$
to obtain $\bm{z}_{H}^{+}$ (with $\bar{\bm{z}}_{H}$ as the initial guess). The minimization can be performed recursively by calling MG/OPT.
6:        Compute the search direction $\bm{e}_{H}=\bm{z}_{H}^{+}-\bar{\bm{z}}_{H}$.
7:        Use a line search to determine $\bm{z}_{h}^{+}=\bar{\bm{z}}_{h}+\alpha\,\bm{I}_{H}^{h}\,\bm{e}_{H}$ satisfying $f_{h}(\bm{z}_{h}^{+})\leq f_{h}(\bar{\bm{z}}_{h})$.
8:     Post-smoothing: Apply $k_{2}$ iterations of the same convergent optimization algorithm to $\min_{\bm{z}}f_{h}(\bm{z})$ to obtain $\bm{z}_{h}^{{(j+1)}}$ (with $\bm{z}_{h}^{+}$ as the initial guess).
9:  end for
Note that the surrogate model optimized at the coarse-grid level is not the original $f_{H}$, but rather
$f_{s}(\bm{z}_{H})=f_{H}(\bm{z}_{H})-\bar{\bm{v}}_{H}^{T}\,\bm{z}_{H}\,,$
where the correction term $-\bar{\bm{v}}_{H}^{T}\bm{z}_{H}$ enforces matching of first-order conditions across levels:
$\nabla f_{h}(\bm{z}_{h})=0\Longrightarrow\bm{I}_{h}^{H}\,\nabla f_{h}(\bm{z}_{h})=0\Longleftrightarrow\nabla f_{s}\left(\bm{I}_{h}^{H}\,\bm{z}_{h}\right)=0.$
The rigorous analysis that validates MG/OPT is given in [13]. Under mild assumptions, the search direction $\bm{e}_{h}$ is a descent direction of $f_{h}$ at $\bar{\bm{z}}_{h}$; that is,
$\nabla f_{h}(\bar{\bm{z}}_{h})^{\top}\bm{e}_{h}<0\,,$
when
$\bm{e}_{H}^{T}\nabla^{2}f_{H}(\bar{\bm{z}}_{H}+\eta\,\bm{e}_{H})\bm{e}_{H}>0\quad\text{for all }0\leq\eta\leq 1\,.$
Notably, the correction term ensures that the algorithm yields a descent direction even when $f_{h}$ and $f_{H}$ are inconsistent. However, the standard MG/OPT framework builds the coarse-grid surrogate $f_{s}$ by matching only the first-order (gradient) information of the fine-grid model. While this generic surrogate guarantees classical multigrid consistency, it neglects higher-order curvature and the strong block structure inherent in ptychography, causing convergence to slow dramatically after only a few iterations. Instead, we propose a more robust coarse-level surrogate, consistent with the fine-level problem, which eliminates the need for a correction term and respects the ptychographic block structure. Section 3 presents this structure-aware surrogate and demonstrates its integration into our multigrid optimization framework.
2 Surrogate Minimization
In this section, we introduce a surrogate objective for Eqn. (1). This surrogate not only majorizes Eqn. (1) (see Section 2.1 for a precise definition of majorization) and iteratively recovers its optimal solution, but is also sufficiently general to encompass the entire PIE family of solvers. An equivalent formulation, presented in [23], has demonstrated superior performance on the general phase-retrieval problem.
Given $\bm{z}^{(j)}$ as the solution at the $j$-th iteration of an optimizer (e.g., rPIE), we define its corresponding quadratic surrogate model as:
$\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$ $=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}P_{k}\bm{z}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}\,,$ (7)
$=\sum_{k=1}^{N}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,.$
where $\mathcal{R}_{k}$ is defined in Definition 1.1. We note that in the definition of $\mathcal{R}_{k}$, the term $\left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}$ is ill-defined if $\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0$ at some index $r$, so that we complement the original definition with the following specification:
$\left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}=\begin{cases}0\,,&\text{if }j=0\,\&\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j-1)}\right)\right)\right]_{r}\,,&\text{if }j\geq 1\,\&\left[\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right]_{r}=0,\\ \left[\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right]_{r}\,,&else\end{cases}$ (8)
The $(j+1)$-th iteration solution is then a numerical solution to:
$\bm{z}^{(j+1)}=\operatornamewithlimits{argmin}_{\bm{z}}\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)\,.$ (9)
Moreover, observing that  Eqn. (7) is expressed as a summation, it is natural to adopt a stochastic optimization strategy. Specifically, in each iteration, we shuffle the indices of the scanning regions $\{k\}_{k=1}^{N}$ and update them sequentially:
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\operatornamewithlimits{argmin}_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\,.$ (10)
One desirable feature of the surrogate model  Eqn. (7) is that it is quadratic. Consequently, an explicit analytical solution is available. To avoid singularity caused by small or even near zero entries of $\bm{Q}$, we further relax $\widetilde{\Phi}_{k}$ by adding a regularization term (similar to the PIE approach):
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\operatornamewithlimits{argmin}_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,.$ (11)
Notice that the update formulation  Eqn. (11) is equivalent to applying one step of the proximal point method [24] to  Eqn. (10).
2.1 Properties of the surrogate model
Our proposed surrogate model in Eqn. (9) enjoys many desirable properties and is designed in accordance with the general principle of majorization [25] (see Figure 2).
Figure 2: Illustration of the majorization property adapted from [26].
Proposition 2.1 (Majorization).
The quadratic surrogate $\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$, as an approximation to the original objective function $\Phi(\bm{z})$ in  Eqn. (1), has the following properties:
  • $\Phi(\bm{z})$ and $\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$ agree at $\bm{z}^{(j)}$:
    $\Phi\left(\bm{z}^{(j)}\right)=\widetilde{\Phi}\left(\bm{z}^{(j)};\bm{z}^{(j)}\right)\,.$
  • $\Phi(\bm{z})$ is dominated by $\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$:
    $\Phi\left(\bm{z}\right)\leq\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)\quad\text{for all}\quad\bm{z}\in\mathbb{C}^{n^{2}}\,.$ (12)
  • The complex gradients of $\Phi(\bm{z})$ and $\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$ agree at $\bm{z}^{(j)}$:
    $\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)=\nabla_{\bm{z}}\widetilde{\Phi}\left(\bm{z}^{(j)};\bm{z}^{(j)}\right)\,.$
The proof of this theorem appears in Appendix B.1. The first statement is straightforward to verify. To show the inequality (12), we deploy the definitions in Eqn. (1) and Eqn. (7) and prove the inequality for every $\Phi_{k}$. The last statement comes from straightforward computation. We call $\widetilde{\Phi}$ a majorization of $\Phi$ due to the inequality Eqn. (12).
Furthermore, minimizing the surrogate (see  Eqn. (9)) provides a descent direction for the original objective function, as implied in the following theorem:
Theorem 2.2 (Convergence and optimality).
Suppose that the quadratic surrogate $\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)$, given in Eqn. (7) has a global minimum $\bm{z}^{(j+1)}$. Then, for all $j=0,1,\dots$, we have
$\Phi\left(\bm{z}^{(j+1)}\right)\leq\Phi\left(\bm{z}^{(j)}\right)\,.$
Additionally,
$\Phi\left(\bm{z}^{(j+1)}\right)=\Phi\left(\bm{z}^{(j)}\right)\iff\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j+1)}\right)=0\,.$
We leave the proof of this theorem to Appendix B.2. The monotonicity inequality follows directly from Proposition 2.1, and the first-order optimality of $\Phi$ follows from the corresponding first-order optimality condition on the quadratic surrogate, with $\mathcal{R}_{k}$ specified in (8)
The convergence rate of the surrogate minimization using  Eqn. (9) is analyzed in the following theorem:
Theorem 2.3 (Convergence rate).
Iterates $\bm{z}^{(j)}$ delivered by the surrogate minimization (see  Eqn. (9)) converge sublinearly with respect to $\Phi$ (see  Eqn. (1)):
$\min_{0\leq j\leq n}\left\|\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)\right\|_{2}^{2}\leq\frac{\left\|\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right\|_{\infty}}{2(n+1)}\Phi\left(\bm{z}^{(0)}\right)\,,$
where $\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\in\mathbb{C}^{n^{2}}$ and $\left\|\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right\|_{\infty}$ is the maximum magnitude among its entries.
We leave the proof of this theorem to Appendix C. The overall derivation follows that in [27], where the increments at each iteration are bounded, and a telescoping-sum argument is applied.
3 MAGPIE
As demonstrated in [20], the multigrid solver improves computational efficiency, while the surrogate majorization model provides favorable convergence properties. We now integrate these approaches into a single, tractable, and robust method, which we term the Multilevel–Adaptive–Guided Ptychographic Iterative Engine (MAGPIE). The name reflects its multigrid structure, adaptive construction of coarse-grid surrogates, and guided fine-grid reconstruction via coarse-grid search directions. We present the algorithm in Section 3.1 and defer discussion of its well-definedness, consistency, descent properties, and automatic regularization selection to Section 3.2.
3.1 Solver development of MAGPIE
In this section, we present the proposed algorithm in a two-level formulation, which can be readily extended to multiple levels.
For the $k$-th scanning region at the $j$-th iteration, we have the fine-grid surrogate model as [^5] To simplify the notation and maintain consistency with earlier definitions, we omit the subscript $h$ from the fine-grid terms.:
$\min_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,,$ (13)
and we define the associated coarse-grid surrogate model as:
$\min_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}_{H}\bm{z}_{H,k}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,.$ (14)
The coarse-grid model represents a weighted average of the fine-grid model, with precise definitions given below:
  • The restriction operator $\bm{I}_{h}^{H}$ and the prolongation operator $\bm{I}_{H}^{h}$ are defined in Eqn. (5) and Eqn. (6) respectively.
  • $\bm{Q}_{H}=\bm{I}_{h}^{H}\bm{Q}\in\mathbb{C}^{(m/2)^{2}}$: downsampled probe by directly applying the restriction operator.
  • $\bm{W}_{\mathcal{R}}\in\mathbb{C}^{m^{2}}$: constant weight vector for downsampling $\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)$.
  • $\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)=\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)$: downsampled revised exit wave.
Both the fine- and coarse-grid subproblems are quadratic and, in principle, admit closed-form solutions. In practice, however, a direct solve is ill-posed: whenever entries of $\bm{Q}$ or $\bm{Q}_{H}$ vanish, the corresponding normal equations lose rank and the solution is no longer unique. As with the full surrogate problem in Eqn. (10), we therefore replace an exact solve by a single proximal-point (relaxation) step with fine-grid regularization term $\bm{u}$ and coarse-grid regularization $\bm{u}_{H}$. Concretely, we define:
  • $\bm{W}_{\bm{z}}\in\mathbb{R}^{m^{2}}_{\geq 0}$: constant weight vector for downsampling the current object estimate $\bm{z}_{k}^{(j)}$.
  • $\bm{z}_{H,k}^{(j)}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}_{k}^{(j)}\right)$: downsampled current object estimate by weighted average pooling.
  • $\bm{W}_{\bm{u}}^{H}\in\mathbb{R}^{(m/2)^{2}}_{\geq 0}$: constant weight vector for downsampling regularization term $\bm{u}$.
  • $\bm{u}_{H}=\bm{W}_{\bm{u}}^{H}\odot\bm{I}_{h}^{H}\bm{u}_{h}$: the coarse-grid regularization term.
First, we derive the update for the coarse-grid model using the proximal point method in Eqn. (14):
$\hat{\bm{z}}_{H,k}^{(j)}=\bm{z}_{H,k}^{(j)}+\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}_{H}\bm{z}_{H,k}^{(j)}\right)\,,$
Correspondingly, the fine-grid update drawn from the coarse level is:
$\widetilde{\bm{z}}_{k}^{(j)}=\bm{z}_{k}^{(j)}+\bm{I}_{H}^{h}\underbrace{\left(\hat{\bm{z}}_{H,k}^{(j)}-\bm{z}_{H,k}^{(j)}\right)}_{\bm{e}^{(j)}_{H,k}}=\bm{z}_{k}^{(j)}+\bm{I}_{H}^{h}\bm{e}^{(j)}_{H,k}\,.$ (15)
Then, we apply the proximal point method on  Eqn. (13) at the fine level with a regularization as:
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\widetilde{\bm{z}}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\widetilde{\bm{z}}_{k}^{(j)}\right)\,.$
For computational purposes, the weights ($\bm{W}_{\bm{z}}$, $\bm{W}_{\mathcal{R}}$, and $\bm{W}_{\bm{u}}$) translate information across levels and must be defined carefully to ensure compatibility with the fine-grid optimization. We present the weight formula in Eqn. (16). These definitions guarantee well-definedness, consistency, descent directions, and stability. We discuss these properties in detail in Section 3.2.
We term this multigrid solver for each surrogate model the Multilevel–Adaptive–Guided Proximal Solver (MAGPS). Specifically, “Adaptive” denotes the automatic construction of coarse-grid surrogate terms using the weights specified in Eqn. (16) without manual tuning, “Guided” emphasizes the directional coarse-to-fine corrections in Eqn. (15), and “Proximal Solver” highlights the use of proximal-point updates within the multigrid framework. The pseudocode for this solver is summarized in Algorithm 3.
The overall approach that addresses ptychographic phase retrieval by iteratively calling MAGPS to minimize the surrogate model over randomly sampled scanning regions is termed Multilevel–Adaptive–Guided Ptychographic Iterative Engine (MAGPIE); its pseudocode is presented in Algorithm 4.
Algorithm 3 $\texttt{MAGPS}\left(\bm{z}^{(j)}_{k},\bm{Q},\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right),\bm{u}\right)$
0:  Object $\bm{z}^{(j)}_{k}$, probe $\bm{Q}$, revised exit wave $\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)$, regularization term $\bm{u}$.
  if current level is coarsest then
     $\bm{z}^{+}_{k}\leftarrow\bm{z}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)$.
  else
     Coarse-Level Preparation:
     Coarse probe: $\bm{Q}_{H}\leftarrow\bm{I}_{h}^{H}\left(\bm{Q}\right)$.
     Set weights for downsampling the object: $\bm{W}_{\bm{z}}\leftarrow\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}$.
     Set weights for downsampling the revised exit wave: $\bm{W}_{\mathcal{R}}\leftarrow\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}$.
     Set weights for coarse-grid regularization: $\bm{W}_{\bm{u}}^{H}\leftarrow\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}$.
     Coarse object: $\bm{z}^{(j)}_{H,k}\leftarrow\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\bm{z}^{(j)}_{k}\right)$.
     Coarse revised exit waves: $\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\leftarrow\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)$.
     Coarse-grid regularization: $\bm{u}_{H}\leftarrow\bm{W}_{\bm{u}}^{H}\odot\bm{I}_{h}^{H}\bm{u}$.
     Recursive Coarse Solve:
     $\widehat{\bm{z}}^{(j)}_{H,k}\leftarrow\texttt{MAGPS}\left(\bm{z}^{(j)}_{H,k},\bm{Q}_{H},\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right),\bm{u}_{H}\right)$.
     Coarse-to-Fine Update:
     Coarse-grid update direction: $\bm{e}^{(j)}_{H,k}\leftarrow\widehat{\bm{z}}^{(j)}_{H,k}-\bm{z}^{(j)}_{H,k}$.
     Update fine-grid estimate: $\widetilde{\bm{z}}^{(j)}_{k}\leftarrow\bm{z}^{(j)}_{k}+\bm{I}_{H}^{h}\bm{e}^{(j)}_{H,k}$.
     Post-Smoothing:
     $\bm{z}^{+}_{k}\leftarrow\widetilde{\bm{z}}^{(j)}_{k}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\widetilde{\bm{z}}^{(j)}_{k}\right)$.
  end if
  return  $\bm{z}^{+}_{k}$
Algorithm 4 $\texttt{MAGPIE}(\bm{z}^{(0)},\bm{Q},\{\bm{d}_{k}\}_{k=1}^{N},\bm{u})$
0:  Object: $\bm{z}^{(0)}$, probe: $\bm{Q}$, measurements: $\{\bm{d}_{k}\}_{k=1}^{N}$, regularization term: $\bm{u}$.
1:  while $\bm{z}^{(j)}$ not converged do
2:     for $k$ randomly drawn from $1$ to $N$ without replacement do
3:        $\bm{z}_{k}^{(j)}\leftarrow\texttt{MAGPS}_{k}\left(\bm{z}^{(j)}_{k},\bm{Q},\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right),\bm{u}\right)$
4:     end for
5:  end while
6:  return  $\bm{z}^{(j)}$
3.2 Converging properties of MAGPIE and weight selection
Different choices of the weight vectors lead to different performance of the algorithm. Now we specify our choices of the weight vectors below. These definitions of $\bm{W}_{\bm{z}}$, $\bm{W}_{\mathcal{R}}$, and $\bm{W}_{\bm{u}}^{H}$ ensure the robust performance of MAGPIE.
$\bm{W}_{\bm{z}}=\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,,\quad\bm{W}_{\mathcal{R}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}=\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot|\bm{Q}|^{2}}{\bm{Q}\odot\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)}\,,\quad\bm{W}_{\bm{u}}^{H}=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$ (16)
As will be shown more clearly in Propositions 3.1, 3.2, and 3.4, and in Theorem 3.3, these specific forms not only prevent singularities, maintain coarse-to-fine consistency, and preserve the descent property, but also ensure stability. It is worth noting that the quantity $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ appears in both $\bm{W}_{\bm{z}}$ and $\bm{W}_{\mathcal{R}}$. Its physical meaning is straightforward: in each $2\times 2$ bin of $|\bm{Q}|^{2}$, $\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}$ computes the average of the four entries and upsamples that average back to the original four positions, whereas $|\bm{Q}|^{2}$ each entry’s original value. As a result,
$\bm{z}_{H,k}^{(j)}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}^{(j)}\right)$
performs a bin-wise weighted average of $\bm{z}_{k}^{(j)}$, assigning greater weight to regions where $\bm{Q}$ has higher magnitude (i.e., higher signal-to-noise ratio).
We first show that all weight vectors are element-wise bounded.
Proposition 3.1 (Well-definedness).
The element-wise magnitudes of the weight vectors defined in Eqn.(16) are bounded. In particular:
$\left\|\bm{W}_{\bm{z}}\right\|_{\infty}\leq 4\,,\quad\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}\leq 4\,,\quad\text{and}\quad\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}\leq 1\,,$
where each weight vector’s infinity norm equals the maximum magnitude among its entries.
To prove this proposition, we realize that $\mathbf{W}_{\mathbf{z}}$ represents the weighted average within a bin and the bin size is $4$. The control for the other two terms comes from Jensen’s inequality used on quadratic functions. The proof is left to Appendix E.
We now show that the coarse-grid surrogate model is consistent with the fine-grid surrogate model.
Proposition 3.2 (Consistency).
Given the weight vectors defined in  Eqn. (16), the coarse-level objective $\widetilde{\Phi}_{H,k}$ (as defined in  Eqn. (14)) is consistent with the fine-level objective $\widetilde{\Phi}_{k}$ (as defined in  Eqn. (13)) in the sense that, for all $\bm{z}_{k}\in\mathbb{C}^{m^{2}}$, given $\bm{z}_{H,k}=\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right)$, we have:
  • $\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)\leq\frac{1}{4}\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}^{2}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$
  • $\left\|\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)\right\|_{2}\leq\frac{1}{2}\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}\,\left\|\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right\|_{2}$
We leave the proof to Appendix F. Most of the proof is a straightforward derivation.
Furthermore, the parameters are chosen so that the coarse-grid solution of Eqn. (14), obtained via the proximal point method, provides a descent direction for the fine-level problem. This is formalized in the following theorem.
Theorem 3.3 (Descent property).
Given the weight vectors defined in  Eqn. (16), a single proximal-point sweep on the coarse surrogate (14), initialized at $\bm{z}^{\prime}_{H,k}=\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\bm{z}^{\prime}_{k})$, provides a descent direction for the fine-grid surrogate (13) at $\bm{z}^{\prime}_{k}$. That is:
$\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\bm{I}_{H}^{h}\left(\bm{e}_{H,k}\right)\leq 0\,,$
where $\bm{I}_{H}^{h}\left(\bm{e}_{H,k}\right)=\bm{I}_{H}^{h}\left(\widehat{\bm{z}_{H,k}}-\bm{z}_{H,k}^{\prime}\right)$, and $\widehat{\bm{z}_{H,k}}$ solves the proximal point on the coarse level:
$\widehat{\bm{z}_{H,k}}=\operatornamewithlimits{argmin}_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{z}_{H,k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}_{H}\cdot\left|\bm{z}_{H,k}-\bm{z}_{H,k}^{\prime}\right|^{2}\,.$ (17)
Moreover,
$\bm{e}_{H,k}$ $=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{z}_{H,k}^{\prime}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right)$ (18)
$=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
The proof of this theorem appears in Appendix G.1. First, we state Proposition G.1 showing a specific form under which the coarse-grid update is a descent direction. Then we show that the coarse-grid update in Eqn. (17) indeed falls into that form.
Whenever a model is represented on both fine and coarse grids, parameters fixed on the fine grid must have consistent counterparts on the coarse grid. We show below that, with $\bm{W}_{\bm{z}},\bm{W}_{\mathcal{R}},\bm{W}_{\bm{u}}^{H}$ designed as in Eqn. (16), the regularization term can be transferred automatically between levels, requiring no manual tuning for the coarse-grid regularization term $\bm{u}_{H}$.
Proposition 3.4 (Automatic regularization selection).
Given a fine-grid regularization term $\bm{u}$ and the weight vector $\bm{W}_{\bm{u}}^{H}$ defined in  Eqn. (16), by choosing
$\bm{u}_{H}=\bm{W}_{\bm{u}}^{H}\bm{I}_{h}^{H}\bm{u}\,,$
the induced coarse-grid regularization term $\bm{u}_{H}$ ensures that, for each scanning region $k$, there exist mappings
$\bm{e}_{k}:\mathbb{C}^{m^{2}}\to\mathbb{C}^{m^{2}}\quad\text{and}\quad\bm{e}_{H,k}:\mathbb{C}^{m^{2}}\to\mathbb{C}^{(m/2)^{2}}\,,$
such that:
  • The mappings $\bm{e}_{k}(\bm{g})$ and $\bm{e}_{H,k}(\bm{g})$ yield the fine-grid update at $\bm{z}_{k}^{\prime}$ and the coarse-grid update at $\bm{z}_{H,k}^{\prime}=\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\bm{z}_{k}^{\prime})$, respectively, for the fine-grid gradient $\bm{g}=\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}(\bm{z}_{k}^{\prime};\,\bm{z}_{k}^{(j)})$.
  • For any coordinate index $r=1,\dots,(m/2)^{2}$, the coarse-grid update satisfies
    $\max_{\|\bm{g}_{1}\|_{\infty}\leq 1}\left|\bm{e}_{H,k}(\bm{g}_{1})\right|_{r}=\max_{\|\bm{g}_{2}\|_{\infty}\leq 1}\left|\bm{I}_{h}^{H}\,\bm{e}_{k}(\bm{g}_{2})\right|_{r}\,.$
    Thus, provided the fine-grid update is stable under a properly tuned regularization term, the maximum magnitude of the coarse-grid update is automatically controlled.
The proof of this proposition appears in Appendix H. The key step uses Eqn. (17), which links the coarse-grid update to the fine-grid surrogate gradient.
4 Numerical Examples
In this section, we present numerical examples to demonstrate the performance of MAGPIE. We compare our solver with classical methods such as rPIE [11] and L-BFGS [28]. For L-BFGS, we set the history size to $5$, which determines the number of recent updates used to approximate the inverse Hessian. To ensure a fair comparison with rPIE, we adopt its regularization term as the fine-grid regularization for MAGPIE, i.e.,
$\bm{u}=\bm{u}^{\texttt{rPIE}}=\alpha\left(\|\bm{Q}\|_{\infty}^{2}-|\bm{Q}|^{2}\right)\,.$
We observe that the computational bottleneck at each iteration stems from the Fourier transforms required to evaluate $\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)$. Since this term is computed only once per scanning region at each iteration, all three algorithms share comparable per-iteration computational costs.
We first describe the problem setup, and then in Sections 4.1, 4.2, and 4.3, we evaluate the performance of the algorithms from three perspectives: the number of multigrid levels, robustness to noise, and the effect of the overlap ratio. Finally, Section 4.4 evaluates the algorithm on a simulated integrated-circuit object to illustrate its practical imaging performance.
Probe:
In our numerical experiments, we use a simulated Fresnel zone plate as the probe (see [29]). The standard probe size is $m=128$. Figure 3 shows the magnitude and phase of the probe. The magnitude displays the characteristic circular structure of a zone plate, while the phase reveals the corresponding alternating concentric patterns. It is worth noting that a large portion of the probe (beyond the center) has a very small magnitude, leading to possible ill-conditioning.
Figure 3: Magnitude (left) and phase (right) of the simulated Fresnel zone plate used as the probe in numerical experiments.
Object:
We test two types of objects. The synthetic object used in our numerical experiments is constructed by combining two standard test images: the Baboon image for magnitude and the Cameraman image for phase. Specifically, the Baboon image is normalized to the range $[0,1]$, representing the object’s magnitude, while the Cameraman image is scaled to span the range $[0,\pi/2]$, providing the object’s phase distribution. The resulting complex-valued object, obtained by multiplying the magnitude by the complex exponential of the phase, exhibits structured patterns and realistic variability in both magnitude and phase. The objects are presented with the resolution of $n=512$. Figure 4 illustrates the magnitude and phase of this synthetic test object.
Figure 4: Magnitude (left) and phase (right) of the synthetic complex-valued object used in numerical experiments. The magnitude is derived from the Baboon image, while the phase is derived from the Cameraman image.
Another synthetic object we use to test our algorithm is more realistic, as shown in Figure 5, which illustrates the magnitude and phase of this test object. This example is inspired by applications in integrated circuit imaging [5].
Figure 5: Magnitude (left) and phase (right) of the realistic complex-valued object used in numerical experiments.
Overlap ratio:
[…]
Noise level:
To mimic realistic photon-counting statistics, we add Poisson noise to every simulated intensity. Given the noiseless value $\widetilde{\bm{d}}_{k}$, the noisy measurement is
$\bm{d}_{k}=\eta\,\mathrm{Poisson}\left(\widetilde{\bm{d}}_{k}/\eta\right)\,,$
where the dimensionless parameter $\eta$ defines the noise level: larger $\eta$ lowers the mean photon count and thus raises the relative noise.
In our experiments, the noise level is set to $\eta=0.05$ by default.
Stopping criterion:
We impose the following stopping criterion for all algorithms:
$\frac{1}{Nm}\sum_{k=1}^{N}\left\|\nabla_{\bm{z}_{k}}\Phi_{k}(\bm{z}_{k})\right\|_{2}<\texttt{tol}\,.$
Quality metrics:
We evaluate reconstruction quality using two metrics: residual and error. For a reconstruction $\bm{z}$, the residual is defined as the value of the loss function $\Phi(\bm{z})$. The error is measured as the absolute $\ell^{2}$-norm difference between the element-wise magnitudes of the reconstruction and the ground truth object, i.e., $\||\bm{z}|-|\bm{z}^{*}|\|_{2}$.
4.1 Effect of multigrid levels on performance
In this experiment, we investigate how the number of coarse levels impacts the performance of MAGPIE. In particular, with $l=1,\cdots,\log_{2}(m)$ levels, we name the corresponding method MAGPIE_$\_l$. Since $m=128$, the maximum level for MAGPIE is $\log_{2}(m)=7$.
In this case, we set the convergence tolerance $\texttt{tol}=10^{-4}$, and the regularization constant $\alpha=0.01$.
The residuals and errors are presented in Figure 6 on the log-log scale. We note that MAGPIE_$\_l$ outperforms rPIE and L-BFGS, and converges faster with each additional coarse level. Therefore, for subsequent tests, we consider only MAGPIE at its deepest level, $l=\log(m)$.
[…]
We showcase the reconstructions obtained with L-BFGS, rPIE, and MAGPIE in Figure 7. The reconstruction produced by MAGPIE exhibits higher quality; in particular, the differences in both magnitude and phase are much smaller and smoother compared to those from L-BFGS and rPIE.
[…]
4.2 Stability under noise
Numerical tests demonstrate that MAGPIE retains high accuracy in the presence of noise. In the following examples, we benchmark it against L-BFGS and rPIE across a range of noise levels. We set the convergence tolerance to $\texttt{tol}=10^{-4}$ and the regularization constant to $\alpha=0.025$ (increasing $\alpha$ for higher noise levels). To assess stability, we add Poisson noise to the measurements at four noise levels: $\eta=0.05,0.1,0.2$, and 0.4.
We present the residuals and errors in Figure 8 and in Figure 9 respectively, on the log-log scale. In all cases, MAGPIE presents better performance in both residuals and errors.
[…]
We showcase the reconstructions of L-BFGS, rPIE, and MAGPIE when noise level $\eta=0.4$ in Figure 10. MAGPIE produces a better reconstruction that has a much smaller and smoother error.
[…]
4.3 Overlap ratio
[…]
4.4 Realistic object
We extend the previous experiments to a realistic object whose entries near the boundary are mostly zero. This test demonstrates that the performance gain of MAGPIE is not solely due to improved reconstruction near the boundary, where fewer scanning regions overlap.
[…]
5 Conclusion
We have presented MAGPIE, a stochastic multigrid minimization framework for ptychographic phase retrieval that accelerates PIE-type updates via rigorously constructed coarse-grid surrogate models. The surrogate guarantees monotonic descent of the original nonconvex objective with sublinear convergence, embeds a multilevel hierarchy in which coarse-grid regularization terms are automatically selected by fine–coarse consistency, and retains the standard rPIE parameter set while significantly improving both convergence speed and reconstruction fidelity.
While we have employed a specific downsampling strategy, alternative approaches are possible. For example, one may define
$\bm{Q}_{H}=\bm{M}_{h}^{H}(\bm{Q})\,,\quad\bm{z}_{H,k}=\bm{I}_{h}^{H}\bm{z}_{k}\,,\quad\text{and }\ \bm{W}_{\mathcal{R}}=\frac{\bm{I}_{H}^{h}\bm{Q}_{H}}{\bm{Q}}\,,$
where $\bm{M}_{h}^{H}$ selects the minimum-norm entry in each bin. In such variants, the properties established in Section 3.2 still hold, although their analysis becomes more challenging.
Looking ahead, we will extend the theoretical framework to blind ptychography and integrate a minibatch strategy for GPU-accelerated real-time feedback. Moreover, the multigrid-surrogate paradigm developed here is readily adaptable to other large-scale phase retrieval and coherent diffractive imaging applications.
6 Acknowledgments
Z.D. was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under Contract DE-AC02-06CH11357. Q.L. acknowledges support from NSF-DMS-2308440, and B.Z. acknowledges support from NSF-DMS-2012292.
Appendix A Computation for Eqn. (2) and Eqn. (4)
We consider the Euclidean space representations $\mathbb{R}^{2m^{2}}$ on the vectors in the complex field $\bm{z}_{k}\in\mathbb{C}^{m^{2}}$ as follows:
$\left\{[\bm{x}_{k}^{\top},\bm{y}_{k}^{\top}]^{\top};\bm{x}_{k}+\bm{y}_{k}i=\bm{z}_{k}\right\}\,.$
With the projection operator:
$\mathcal{R}_{k}(\bm{z}_{k})=\mathcal{F}^{-1}\left[\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta(\mathcal{F}(\bm{Q}\odot\bm{z}_{k}))\right)\right]\,,$
we define the objective function in the $k$-th scanning region:
$\Phi_{k}(\bm{z}_{k})=\frac{1}{2}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right\|_{2}^{2}\,,$
and find its complex gradient:
$\nabla_{\bm{z}_{k}}\Phi_{k}=\nabla_{\bm{x}_{k}}\Phi_{k}+i\nabla_{\bm{y}_{k}}\Phi_{k}\,.$
Firstly, the Parseval’s theorem for the discrete Fourier transform gives:
$\|\mathcal{F}(\bm{z}_{k})\|_{2}^{2}=m^{2}\|\bm{z}_{k}\|_{2}^{2}\quad\text{and}\quad\|\mathcal{F}^{-1}(\bm{z}_{k})\|_{2}^{2}=\frac{1}{m^{2}}\|\bm{z}_{k}\|_{2}^{2}\,.$
Consequently, for all $\bm{v}\in\mathbb{C}^{m^{2}}$ we have
$\langle\mathcal{F}(\bm{v}),\mathcal{F}(\bm{v})\rangle=\langle\bm{v},\mathcal{F}^{\ast}\mathcal{F}(\bm{v})\rangle=m^{2}\langle\bm{v},\bm{v}\rangle\,,$
which implies
$\frac{1}{m^{2}}\mathcal{F}^{\ast}\mathcal{F}(\bm{v})=\bm{I}\implies\mathcal{F}^{-1}=\frac{1}{m^{2}}\mathcal{F}^{\ast}\,.$
A.1 Computation for Eqn. (2)
We find
$\nabla_{\bm{z}_{k}}\Phi_{k}=\overline{\bm{Q}}\odot(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k}))\,.$
For the $k$-th scanning region, we have
$\Phi_{k}(\bm{z}_{k})$ $=\frac{1}{2}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right\|_{2}^{2}\,,$
$=\frac{1}{2m^{2}}\left\|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})-\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta(\mathcal{F}(\bm{Q}\odot\bm{z}_{k}))\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2m^{2}}\left\||\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|-\sqrt{\bm{d}_{k}}\right\|_{2}^{2}\,.$
Using the chain rule for the cogradient operator defined in Chapter 4.2 of [22], we have
$\frac{\partial\,|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|}{\partial\,\bm{z}_{k}}$ $=\frac{\partial\,|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|}{\partial\,\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}\frac{\partial\,\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}{\partial\,\bm{z}_{k}}+\frac{\partial\,|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|}{\partial\,\overline{\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}}\frac{\partial\,\overline{\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}}{\partial\,\bm{z}_{k}}\,,$
$=\operatorname{diag}\left(\frac{\overline{\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}}{2|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|}\right)\mathcal{F}\operatorname{diag}(\bm{Q})\,.$
We note that $\frac{\partial\,\overline{\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}}{\partial\,\bm{z}_{k}}$ can be computed as 0.
Consequently, we have
$\frac{\partial\,\Phi_{k}}{\partial\,\bm{z}_{k}}$ $=\frac{1}{2m^{2}}\frac{\partial\,\left\||\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|-\sqrt{\bm{d}_{k}}\right\|_{2}^{2}}{\partial\,\bm{z}_{k}}\,,$
$=\frac{1}{m^{2}}\left(|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|-\sqrt{\bm{d}_{k}}\right)\frac{\partial\,|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|}{\partial\,\bm{z}_{k}}\,,$
$=\frac{1}{m^{2}}\left(|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|-\sqrt{\bm{d}_{k}}\right)\operatorname{diag}\left(\frac{\overline{\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}}{2|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|}\right)\mathcal{F}\operatorname{diag}(\bm{Q})\,.$
Finally, by definition, we obtain
$\nabla_{\bm{z}_{k}}\Phi_{k}$ $=\nabla_{\bm{x}_{k}}\Phi_{k}+i\nabla_{\bm{y}_{k}}\Phi_{k}\,,$
$=\left(\frac{\partial\,\Phi_{k}}{\partial\,\bm{x}_{k}}+i\frac{\partial\,\Phi_{k}}{\partial\,\bm{y}_{k}}\right)^{\top}\,,$
$=2\frac{\partial\,\Phi_{k}}{\partial\,\bm{z}_{k}}^{*}\,,$
$=\operatorname{diag}\left(\overline{\bm{Q}}\right)\left(\frac{1}{m^{2}}\mathcal{F}^{*}\right)\left[\operatorname{diag}\left(\frac{\mathcal{F}(\bm{Q}\odot\bm{z}_{k})}{|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|}\right)\left(|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|-\sqrt{\bm{d}_{k}}\right)\right]\,,$
$=\overline{\bm{Q}}\odot\mathcal{F}^{-1}\left[\exp\left(i\theta(\mathcal{F}(\bm{Q}\odot\bm{z}_{k}))\right)\odot\left(|\mathcal{F}(\bm{Q}\odot\bm{z}_{k})|-\sqrt{\bm{d}_{k}}\right)\right]\,,$
$=\overline{\bm{Q}}\odot\mathcal{F}^{-1}\left[\mathcal{F}(\bm{Q}\odot\bm{z}_{k})-\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta(\mathcal{F}(\bm{Q}\odot\bm{z}_{k}))\right)\right]\,,$
$=\overline{\bm{Q}}\odot\left[\bm{Q}\odot\bm{z}_{k}-\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\odot\exp\left(i\theta(\mathcal{F}(\bm{Q}\odot\bm{z}_{k}))\right)\right)\right]\,,$
$=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right)\,.$
Note that this computation does not work when $\mathcal{F}(\bm{Q}\odot\bm{z}_{k}))_{r}=0$ at some index $r$. This issue can be prevented by finding the gradient using the third statement of Proposition 2.1.
A.2 Computation for Eqn. (4)
We find
$\nabla_{\bm{z}_{k}}\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\odot\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\,,$
where
$\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,.$
For the first term, we have
$\frac{\partial\,\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}}{\partial\,\bm{z}_{k}}$ $=\frac{\partial\,\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\cdot\overline{\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)}}{\partial\,\bm{z}_{k}}$
$=\overline{\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)}\operatorname{diag}(\bm{Q})\,.$
For the second term,
$\frac{\partial\,\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}}{\partial\,\bm{z}_{k}}$ $=\frac{\partial\,\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\cdot\left(\bm{u}\odot\overline{\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)}\right)}{\partial\,\bm{z}_{k}}$
$=\bm{u}\odot\overline{\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)}\,.$
Therefore,
$\nabla_{\bm{z}_{k}}\Phi^{\texttt{PIE}}_{k}$ $=2\frac{\partial\,\Phi^{\texttt{PIE}}_{k}}{\partial\,\bm{z}_{k}}^{*}\,,$
$=2\cdot\frac{1}{2}\cdot\left(\operatorname{diag}\left(\overline{\bm{Q}}\right)\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\odot\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\right)\,,$
$=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\odot\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\,.$
Appendix B Proof for Proposition 2.1, and Theorem 2.2
We begin with a lemma and a proposition.
Lemma B.1.
For any complex numbers $x,y\in\mathbb{C}$ and $d\geq 0$,
$\left|x-de^{i\theta(x)}\right|^{2}\leq\left|x-de^{i\theta(y)}\right|^{2}\,.$
Additionally, equality implies either $d=0$, $x=0$, or $\theta(x)=\theta(y)$.
Proof.
The first term can be expanded as:
$\left|x-de^{i\theta(x)}\right|^{2}$ $=\left||x|e^{i\theta(x)}-de^{i\theta(x)}\right|^{2}\,,$
$=\left||x|-d\right|^{2}\left|e^{i\theta(x)}\right|^{2}\,,$
$=|x|^{2}+d^{2}-2|x|d\,.$
The second term can be expanded as:
$\left|x-de^{i\theta(y)}\right|^{2}$ $=\left||x|e^{i\theta(x)}-de^{i\theta(y)}\right|^{2}\,,$
$=\left(|x|e^{i\theta(x)}-de^{i\theta(y)}\right)\left(|x|e^{-i\theta(x)}-de^{-i\theta(y)}\right)\,,$
$=|x|^{2}+d^{2}-|x|d\left(e^{i(\theta(x)-\theta(y))}+e^{i(\theta(y)-\theta(x))}\right)\,,$
$=|x|^{2}+d^{2}-2|x|d\cos(\theta(x)-\theta(y))$
The inequality follows from the fact that $\cos(\theta(x)-\theta(y))\leq 1$ and that $d\geq 0$. Additionally, equality implies either $d=0$, $|x|=0$, or $\cos(\theta(x)-\theta(y))$, which is equivalent to either $d=0$, $x=0$, or $\theta(x)=\theta(y)$. ∎
Proposition B.2.
For each scanning region indexed by $k=1,\dots,N$,
$\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}\right)\right\|_{2}^{2}\leq\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,.$
Additionally, equality implies $\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)$ except at the zero entries of $\bm{d}_{k}$ or at the zero entries of $\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)$.
Proof.
Using Parseval’s theorem for the discrete Fourier transform, it suffices to show:
$\left\|\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)-\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)\right)\right)\right\|_{2}^{2}\leq\left\|\mathcal{F}\left(\bm{Q}\bm{z}_{k}\right)-\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}^{2}\,.$
The inequality and the additional equality outcomes follow from Lemma B.1, which applies to the inequality element-wise. ∎
B.1 Proof for Proposition 2.1
Proof.
It is straightforward to check the first statement.
For the second statement, it follows from Proposition B.2 that $\Phi_{k}\left(\bm{z}_{k}\right)\leq\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ for each scanning region. Hence,
$\Phi\left(\bm{z}\right)=\sum_{k=1}^{N}\Phi_{k}\left(\bm{z}_{k}\right)\leq\sum_{k=1}^{N}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)\,.$
For the last statement, we consider equivalent representations of the two objective functions using the real and imaginary parts of the input $\bm{z}=\bm{x}+i\bm{y}$:
$\Phi\left(\bm{z}\right)\equiv\Phi\left(\bm{x},\bm{y}\right)\quad\text{and}\quad\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)\equiv\widetilde{\Phi}\left(\bm{x},\bm{y};\bm{z}^{(j)}\right)\,.$
With $\bm{z}^{(j)}=\bm{x}^{(j)}+i\bm{y}^{(j)}$, it suffices to show:
$\nabla_{\bm{x}}\Phi\left(\bm{x}^{(j)},\bm{y}^{(j)}\right)=\nabla_{\bm{x}}\widetilde{\Phi}\left(\bm{x}^{(j)},\bm{y}^{(j)};\bm{z}^{(j)}\right)\quad\text{and}\quad\nabla_{\bm{y}}\Phi\left(\bm{x}^{(j)},\bm{y}^{(j)}\right)=\nabla_{\bm{y}}\widetilde{\Phi}\left(\bm{x}^{(j)},\bm{y}^{(j)};\bm{z}^{(j)}\right)\,.$
Bythe previous two statements and definition of the gradient, for a small perturbation $\epsilon\bm{h}$ such that $\epsilon>0$ and $\bm{h}\in\mathbb{R}^{m^{2}}$, we have
$\epsilon\left\langle\nabla_{\bm{x}}\Phi\left(\bm{x}^{(j)},\bm{y}^{(j)}\right),\bm{h}\right\rangle+o(\epsilon)$ $=\Phi\left(\bm{x}^{(j)}+\epsilon\bm{h},\bm{y}^{(j)}\right)-\Phi\left(\bm{x}^{(j)},\bm{y}^{(j)}\right)\,,$
$\leq\widetilde{\Phi}\left(\bm{x}^{(j)}+\epsilon\bm{h},\bm{y}^{(j)};\bm{z}^{(j)}\right)-\widetilde{\Phi}\left(\bm{x}^{(j)},\bm{y}^{(j)};\bm{z}^{(j)}\right)$
$=\epsilon\left\langle\nabla_{\bm{x}}\widetilde{\Phi}\left(\bm{x}^{(j)},\bm{y}^{(j)};\bm{z}^{(j)}\right),\bm{h}\right\rangle+o(\epsilon)\,.$
Noting that $\epsilon>0$ and $\bm{h}\in\mathbb{R}^{m^{2}}$ are arbitrary, we have
$\nabla_{\bm{x}}\Phi\left(\bm{x}^{(j)},\bm{y}^{(j)}\right)=\nabla_{\bm{x}}\widetilde{\Phi}\left(\bm{x}^{(j)},\bm{y}^{(j)};\bm{z}^{(j)}\right)\,.$
Similarly, we can show the equality for $\nabla_{\bm{y}}$ and conclude that
$\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)=\nabla_{\bm{z}}\widetilde{\Phi}\left(\bm{z}^{(j)};\bm{z}^{(j)}\right)\,.$
B.2 Proof for Theorem 2.2
Proof.
We separate the proof into two parts:
  1. 1.
    Inequality: Since we obtain $\bm{z}^{(k+1)}$ by solving
    $\bm{z}^{(k+1)}=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,,$
    we have
    $\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\leq\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}=\Phi\left(\bm{z}^{(j)}\right)\,.$ (19)
    Then, by Proposition B.2, we have
    $\Phi\left(\bm{z}^{(j+1)}\right)=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)\right\|_{2}^{2}\leq\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{(j+1)}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,.$ (20)
    Combining Equation Eqn. (19) and Equation Eqn. (20) completes the proof of the inequality.
  2. 2.
    Optimality: By Proposition B.2, the inequality Eqn. (20) achieves equality only when $\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)$ except at the zero entries of $\bm{d}_{k}$ or at the zero entries of $\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)$ for all $k=1,2,\dots,N$.
    Since at the zero entries $\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)$, we impose $\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)\right)=\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)$ in the Eqn. (8), we have, for all $k=1,2,\dots,N$,
    $\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)$ $=\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j+1)}\right)\right)\right)\right)\,,$
    $=\mathcal{F}^{-1}\left(\sqrt{\bm{d}_{k}}\exp\left(i\theta\left(\mathcal{F}\left(\bm{Q}\bm{z}_{k}^{(j)}\right)\right)\right)\right)\,,$
    $=\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\,.$
    That leads to
    $\bm{z}^{(j+1)}$ $=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j)}_{k}\right)\right\|_{2}^{2}\,,$
    $=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}^{\prime}_{k}-\mathcal{R}_{k}\left(\bm{z}^{(j+1)}_{k}\right)\right\|_{2}^{2}\,,$
    $=\operatornamewithlimits{argmin}_{\bm{z}^{\prime}}\widetilde{\Phi}\left(\bm{z}^{\prime};\bm{z}^{(j+1)}\right)\,.$
    By the optimality of a quadratic objective function, we have
    $\nabla_{\bm{z}}\widetilde{\Phi}\left(\bm{z}^{(j+1)};\bm{z}^{(j+1)}\right)=\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}^{(j+1)}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j+1)}\right)\right)=0\,.$
    Therefore, we can conclude that
    $\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j+1)}\right)=\nabla_{\bm{z}}\widetilde{\Phi}\left(\bm{z}^{(j+1)};\bm{z}^{(j+1)}\right)=0\,.$
Appendix C Proof for Theorem 2.3
Proof.
By Proposition 2.1, Equation Eqn. (9), and Hölder’s inequality, we have
$\Phi\left(\bm{z}^{(j+1)}\right)-\Phi\left(\bm{z}^{(j)}\right)$ $\leq\widetilde{\Phi}\left(\bm{z}^{(j+1)};\bm{z}^{(j)}\right)-\widetilde{\Phi}\left(\bm{z}^{(j)};\bm{z}^{(j)}\right)\,,$
$\leq\widetilde{\Phi}\left(\bm{z};\bm{z}^{(j)}\right)-\widetilde{\Phi}\left(\bm{z}^{(j)};\bm{z}^{(j)}\right)\,,$
$=\frac{1}{2}\sum_{k=1}^{N}\left[\left\|\bm{Q}P_{k}\bm{z}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}-\left\|\bm{Q}P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right\|_{2}^{2}\right]\,,$
$=\frac{1}{2}\sum_{k=1}^{N}\bigg{[}\left\|\bm{Q}P_{k}\bm{z}\right\|^{2}-\overline{\bm{Q}P_{k}\bm{z}}\cdot\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)-\bm{Q}P_{k}\bm{z}\cdot\overline{\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}$
$\quad\quad\quad-\left\|\bm{Q}P_{k}\bm{z}^{(j)}\right\|^{2}+\overline{\bm{Q}P_{k}\bm{z}^{(j)}}\cdot\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)+\bm{Q}P_{k}\bm{z}^{(j)}\cdot\overline{\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}\bigg{]}\,,$
$=\frac{1}{2}\sum_{k=1}^{N}\bigg{[}\left\|\bm{Q}P_{k}\bm{z}-\bm{Q}P_{k}\bm{z}^{(j)}\right\|^{2}+\overline{\bm{Q}P_{k}\bm{z}-\bm{Q}P_{k}\bm{z}^{(j)}}\cdot\left(\bm{Q}P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)\right)$
$\quad\quad\quad+\left(\bm{Q}P_{k}\bm{z}-\bm{Q}P_{k}\bm{z}^{(j)}\right)\cdot\overline{\bm{Q}P_{k}\bm{z}^{(j)}-\mathcal{R}_{k}\left(P_{k}\bm{z}^{(j)}\right)}\bigg{]}\,,$
$=\frac{1}{2}\sum_{k=1}^{N}\bigg{[}\left(P_{k}^{\top}|\bm{Q}|^{2}\right)\cdot\left|\bm{z}-\bm{z}^{(j)}\right|^{2}+\overline{\bm{z}-\bm{z}^{(j)}}\cdot P_{k}^{\top}\nabla_{\bm{z}_{k}}\Phi_{k}\left(\bm{z}_{k}^{(j)}\right)$
$\quad\quad\quad\quad+\left(\bm{z}-\bm{z}^{(j)}\right)\cdot\overline{P_{k}^{\top}\nabla_{\bm{z}_{k}}\Phi_{k}\left(\bm{z}_{k}^{(j)}\right)}\bigg{]}\,,$
$=\left(\frac{1}{2}\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right)\cdot\left|\bm{z}-\bm{z}^{(j)}\right|^{2}+\overline{\bm{z}-\bm{z}^{(j)}}\cdot\left(\frac{1}{2}\sum_{k=1}^{N}P_{k}^{\top}\nabla_{\bm{z}_{k}}\Phi_{k}\left(\bm{z}_{k}^{(j)}\right)\right)$
$\quad\quad\quad+\left(\bm{z}-\bm{z}^{(j)}\right)\cdot\overline{\frac{1}{2}\sum_{k=1}^{N}P_{k}^{\top}\nabla_{\bm{z}_{k}}\Phi_{k}\left(\bm{z}_{k}^{(j)}\right)}\,,$
$\leq\frac{1}{2}\left\|\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right\|_{\infty}\left\|\bm{z}-\bm{z}^{(j)}\right\|_{2}^{2}+\overline{\bm{z}-\bm{z}^{(j)}}\cdot\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)$
$\quad\quad\quad+\left(\bm{z}-\bm{z}^{(j)}\right)\cdot\overline{\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)}\,.$
We denote $\frac{1}{2}\left\|\sum_{k=1}^{N}P_{k}^{\top}|\bm{Q}|^{2}\right\|_{\infty}$ by $L$. Then, the choice $\bm{z}=\bm{z}^{(j)}-\frac{1}{L}\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)$ yields:
$\Phi\left(\bm{z}^{(j)}\right)-\Phi\left(\bm{z}^{(j+1)}\right)\geq\frac{1}{L}\left\|\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)\right\|_{2}^{2}\,.$
A telescoping sum gives:
$\frac{n+1}{L}\min_{0\leq j\leq n}\left\|\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)\right\|_{2}^{2}$ $\leq\frac{1}{L}\sum_{j=0}^{n}\left\|\nabla_{\bm{z}}\Phi\left(\bm{z}^{(j)}\right)\right\|_{2}^{2}\,,$
$\leq\Phi\left(\bm{z}^{(0)}\right)-\Phi\left(\bm{z}^{(n+1)}\right)\,,$
$\leq\Phi\left(\bm{z}^{(0)}\right)\,.$
Appendix D Useful Lemmas
Lemma D.1.
For any $\bm{A}\in\mathbb{C}^{n_{h}^{2}}$ and $\bm{B}\in\mathbb{C}^{n_{H}^{2}}$, we have
$\bm{B}\odot\bm{I}_{h}^{H}\bm{A}=\bm{I}_{h}^{H}((\bm{I}_{H}^{h}\bm{B})\odot\bm{A})\,.$
Note that when $\bm{A}=\bm{J}$, a vector of ones, we have
$\bm{B}=\bm{I}_{h}^{H}\bm{I}_{H}^{h}\bm{B}\implies\bm{I}_{h}^{H}\bm{I}_{H}^{h}=\bm{I},\text{ the identity matrix}\,.$
Proof.
Note that using the binning strategy, each column of $\bm{I}_{h}^{H}$ is non-zero only at one entry, i.e.,
$(\bm{I}_{h}^{H})_{ij}(\bm{I}_{h}^{H})_{kj}=\frac{1}{4}(\bm{I}_{h}^{H})_{ij}\delta_{ik}\,,$
and by definition, we have
$\bm{I}_{H}^{h}=4(\bm{I}_{h}^{H})^{\top}\,.$
For all entries in the LHS and RHS, we have
$\texttt{LHS}_{i}=\bm{A}_{i}(\bm{I}_{h}^{H}\bm{B})_{i}=\bm{A}_{i}\sum_{j}(\bm{I}_{h}^{H})_{ij}\bm{B}_{j}\,.$
$\texttt{RHS}_{i}$ $=(\bm{I}_{h}^{H}[(\bm{I}_{H}^{h}\bm{A})\odot\bm{B}])_{i}\,,$
$=\sum_{j}(\bm{I}_{h}^{H})_{ij}[(\bm{I}_{H}^{h}\bm{A})\odot\bm{B}]_{j}\,,$
$=\sum_{j}(\bm{I}_{h}^{H})_{ij}\left[\sum_{k}(\bm{I}_{H}^{h})_{jk}\bm{A}_{k}\right]\bm{B}_{j}\,,$
$=\sum_{k}\bm{A}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}(\bm{I}_{H}^{h})_{jk}\bm{B}_{j}\,,$
$=\sum_{k}\bm{A}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}4(\bm{I}_{h}^{H})_{kj}\bm{B}_{j}\,,$
$=\sum_{k}\bm{A}_{k}\sum_{j}(\bm{I}_{h}^{H})_{ij}\delta_{ik}\bm{B}_{j}\,,$
$=\bm{A}_{i}\sum_{j}(\bm{I}_{h}^{H})_{ij}\bm{B}_{j}\,.$
Lemma D.2.
For any $\bm{B}\in\mathbb{C}^{n_{H}^{2}}$ and any function $f:\mathbb{C}\to\mathbb{C}$, we define the operators
$\bm{f}_{H}:\mathbb{C}^{n_{H}^{2}}\to\mathbb{C}^{n_{H}^{2}}\quad\text{and}\quad\bm{f}_{h}:\mathbb{C}^{n_{h}^{2}}\to\mathbb{C}^{n_{h}^{2}}\,,$
to be the element-wise applications of $f$, i.e., for any input vector the operator applies $f$ to every entry. Then, for any $\bm{C}\in\mathbb{C}^{n_{H}^{2}}$,
$\bm{f}_{h}\left(\bm{I}_{H}^{h}\bm{C}\right)=\bm{I}_{H}^{h}\bm{f}_{H}(\bm{C})\,.$
Lemma D.3.
For any $\bm{B}$, and $\bm{C}\in\mathbb{C}^{n_{H}^{2}}$ and any binary function $g:\mathbb{C}\otimes\mathbb{C}\to\mathbb{C}$, we define the operators
$\bm{g}_{H}:\mathbb{C}^{n_{H}^{2}}\otimes\mathbb{C}^{n_{H}^{2}}\to\mathbb{C}^{n_{H}^{2}}\quad\text{and}\quad\bm{g}_{h}:\mathbb{C}^{n_{h}^{2}}\otimes\mathbb{C}^{n_{h}^{2}}\to\mathbb{C}^{n_{h}^{2}}\,,$
to be the element-wise applications of $g$. Then, for any $\bm{B},\bm{C}\in\mathbb{C}^{n_{H}^{2}}$,
$\bm{g}_{h}\left(\bm{I}_{H}^{h}\bm{B},\bm{I}_{H}^{h}\bm{C}\right)=\bm{I}_{H}^{h}\bm{g}_{H}(\bm{B},\bm{C})\,.$
Lemma D.2 and Lemma D.3 are straightforward to check by comparing corresponding entries of both sides, so we omit the proofs.
Lemma D.4.
The operator norms of $\bm{I}_{h}^{H}$ and $\bm{I}_{H}^{h}$ induced by the 2-norm are:
$\left\|\bm{I}_{h}^{H}\right\|_{2}=\frac{1}{2}\quad\text{and}\quad\left\|\bm{I}_{H}^{h}\right\|_{2}=2\,.$
The operator norms of $\bm{I}_{h}^{H}$ and $\bm{I}_{H}^{h}$ induced by the $\infty$-norm are:
$\left\|\bm{I}_{h}^{H}\right\|_{\infty}=1\quad\text{and}\quad\left\|\bm{I}_{H}^{h}\right\|_{\infty}=1\,.$
Proof.
For any vector $\bm{v}\in\mathbb{C}^{n_{H}^{2}}$, the prolongation operator $\bm{I}_{H}^{h}$ interpolates $\bm{v}$ by replicating each entry uniformly across its corresponding bin. Hence,
$\left\|\bm{I}_{H}^{h}\bm{v}\right\|_{2}=2\left\|\bm{v}\right\|_{2}\quad\text{and}\quad\left\|\bm{I}_{H}^{h}\bm{v}\right\|_{\infty}=\left\|\bm{v}\right\|_{\infty}\,.$
This leads to
$\left\|\bm{I}_{H}^{h}\right\|_{2}=2\quad\text{and}\quad\left\|\bm{I}_{H}^{h}\right\|_{\infty}=1\,.$
For any vector $\bm{w}\in\mathbb{C}^{n_{h}^{2}}$, the restriction operator $\bm{I}_{h}^{H}$ downsamples $\bm{w}$ by taking average in bins of 4 entries. Let $\mathcal{B}_{j}$ denote the set of indices of the entries in the bin corresponding to the entry $\left(\bm{I}_{h}^{H}\bm{w}\right)_{j}$. Then, by Cauchy–Schwarz inequality,
$\left|\left(\bm{I}_{h}^{H}\bm{w}\right)_{j}\right|=\left|\frac{1}{4}\sum_{i\in\mathcal{B}_{j}}\bm{w}_{i}\right|\leq\frac{1}{2}\sqrt{\sum_{i\in\mathcal{B}_{j}}|\bm{w}_{i}|^{2}}\,.$
Hence, we have
$\left\|\bm{I}_{h}^{H}\bm{w}\right\|_{2}^{2}=\sum_{j=1}^{n_{H}^{2}}\left|\left(\bm{I}_{h}^{H}\bm{w}\right)_{j}\right|^{2}\leq\frac{1}{4}\sum_{j=1}^{n_{H}^{2}}\sum_{i\in\mathcal{B}_{j}}|\bm{w}_{i}|^{2}=\frac{1}{4}\left\|\bm{w}\right\|_{2}^{2}\,.$
Notice that equality is achieved when $\bm{w}$ is a vector of ones, by definition, we have
$\left\|\bm{I}_{h}^{H}\right\|_{2}=\frac{1}{2}\,.$
Similarly, by triangle inequality,
$\left|\left(\bm{I}_{h}^{H}\bm{w}\right)_{j}\right|=\left|\frac{1}{4}\sum_{i\in\mathcal{B}_{j}}\bm{w}_{i}\right|\leq\frac{1}{4}\sum_{i\in\mathcal{B}_{j}}|\bm{w}_{i}|\leq\max_{i\in\mathcal{B}_{j}}|\bm{w}_{i}|\,,$
which leads to
$\left\|\bm{I}_{h}^{H}\bm{w}\right\|_{\infty}=\max_{j=1,\cdots,n_{H}^{2}}\left|\left(\bm{I}_{h}^{H}\bm{w}\right)_{j}\right|\leq\max_{i=1,\cdot,n_{h}^{2}}|\bm{w}_{i}|=\left\|\bm{w}\right\|_{\infty}\,.$
The equality is also achieved when $\bm{w}$ is a vector of ones. Hence, we have
$\left\|\bm{I}_{h}^{H}\right\|_{\infty}=1\,.$
Appendix E Proof for Proposition 3.1
Proof.
We provide proofs to the statements in the same order.
  • By Lemma D.1 and Lemma D.2, we have
    $\bm{I}_{h}^{H}{\bm{W}_{\bm{z}}}$ $=\bm{I}_{h}^{H}\left(\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\,,$
    $=\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)|\bm{Q}|^{2}\right)\,,$
    $=\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\bm{I}_{h}^{H}\left(|\bm{Q}|^{2}\right)\,,$
    $=\bm{J}\,,\quad\text{a vector of ones}\,.$
    For the sake of a contradiction, we assume that $(\bm{W}_{\bm{z}})_{r}>4$ at some index $r$. Since $\bm{W}_{\bm{z}}\geq 0$, we have $\bm{I}_{h}^{H}{\bm{W}_{\bm{z}}}\neq\bm{J}$ because in the bin of index $r$, the average is greater than 1. That is a contradiction to the first statement. Therefore, $\bm{W}_{\bm{z}}$ element-wise bounded by $4$.
  • First, we find that
    $\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|=\frac{|(\bm{I}_{H}^{h}\bm{Q}_{H})||\bm{Q}|^{2}}{|\bm{Q}|\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}\right)}=\frac{\left(\bm{I}_{H}^{h}|\bm{Q}_{H}|\right)|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\leq\frac{\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|\right)|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$
    Then, by Lemma D.1, Lemma D.2, and Lemma D.3, we have
    $\bm{I}_{h}^{H}\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|$ $\leq\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|\right)|\bm{Q}|}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\,,$
    $=\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{\bm{I}_{h}^{H}|\bm{Q}|}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)|\bm{Q}|\right)\,,$
    $=\left(\frac{\bm{I}_{h}^{H}|\bm{Q}|}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\bm{I}_{h}^{H}|\bm{Q}|\,,$
    $=\frac{(\bm{I}_{h}^{H}|\bm{Q}|)^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\leq 1\,,$
    where the last element-wise inequality follows from Jensen’s inequality applied to each bin. Hence, by exactly the same argument as the proof of the previous statement, we have
    $\bm{I}_{h}^{H}\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|\leq 1\implies\left|\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right|\leq 4\,,$
    where the inequalities are satisfied element-wise. We note the sharp bound for the last inequality is $3/2$, which is much harder to find.
  • By applying Jensen’s inequality to each bin,
    $|\bm{Q}_{H}|^{2}\leq\left(\bm{I}_{h}^{H}|\bm{Q}|\right)^{2}\leq\bm{I}_{h}^{H}|\bm{Q}|^{2}\implies\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\leq 1\,.$
Appendix F Proof for Proposition 3.2
Proof.
By definition, we have
$\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)$ $=\frac{1}{2}\left\|\bm{Q}_{H}\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left((\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\bm{Q}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\bm{Q}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}^{2}\,,$
$=\frac{1}{2}\left\|\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}^{2}\,,$
$\leq\frac{1}{2}\left\|\bm{I}_{h}^{H}\right\|_{2}^{2}\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}^{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}\,,$
$=\frac{\left\|\bm{W}_{\mathcal{R}}\right\|_{\infty}^{2}}{4}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\,.$
where the second equality follows from Lemma D.1, and the last inequality follows from Lemma D.4.
Similarly, we have
$\left\|\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)\right\|_{2}$ $=\left\|\overline{\bm{Q}_{H}}\left(\bm{Q}_{H}\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right)-\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,,$
$=\left\|\overline{\bm{Q}_{H}}\bm{I}_{h}^{H}\left(\bm{W}_{\mathcal{R}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,,$
$=\left\|\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\right\|_{2}\,,$
$=\left\|\bm{I}_{h}^{H}\left(\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right)\right\|_{2}\,,$
$\leq\left\|\bm{I}_{h}^{H}\right\|_{2}\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}\left\|\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)\right\|_{2}\,.$
Since $\left\|\bm{I}_{h}^{H}\right\|_{2}=\frac{1}{2}$ by Lemma D.4, it suffices to check $\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}=\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}$ and it is bounded by 1. To this end, by Lemma D.2, Lemma D.3, we find
$\left\|\frac{\left(\bm{I}_{H}^{h}\overline{\bm{Q}_{H}}\right)\bm{W}_{\mathcal{R}}}{\overline{\bm{Q}}}\right\|_{\infty}$ $=\left\|\frac{\overline{\bm{I}_{H}^{h}\bm{Q}_{H}}}{\overline{\bm{Q}}}\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\right\|_{\infty}\,,$
$=\left\|\frac{|\bm{I}_{H}^{h}\bm{Q}_{H}|^{2}}{|\bm{Q}|^{2}}\frac{|\bm{Q}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right\|_{\infty}\,,$
$=\left\|\frac{\bm{I}_{H}^{h}|\bm{Q}_{H}|^{2}}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right\|_{\infty}\,,$
$=\left\|\bm{I}_{H}^{h}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right\|_{\infty}\,,$
$=\left\|\bm{W}_{\bm{u}}^{H}\right\|_{\infty}\,.$
where the last equality is discussed in the proof of Lemma D.4. Sharpness of the second inequality is straightforward to check by setting both $\bm{Q}$ and $\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ as vectors of ones, in which case $\nabla_{\bm{z}_{H,k}}\widetilde{\Phi}_{H,k}\left(\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\bm{z}_{k}\right);\bm{z}_{k}^{(j)}\right)$ is also a vector of ones. ∎
Appendix G Proof for Theorem 3.3
Proposition G.1.
Suppose that
$\bm{e}_{H,k}=-\bm{A}\odot\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
for some nonnegative vector $\bm{A}\in\mathbb{R}^{m_{H}^{2}}$. Then, $\bm{I}_{H}^{h}\left(\bm{e}_{H,k}\right)$ is a descent direction, i.e.
$\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{e}_{H,k}\right)\leq 0\,.$
Additionally, if $\bm{A}$ is positive, then equality implies $\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)=0$.
Proof.
Since $\operatorname{diag}(\bm{A})$ is positive semidefinite, we have
$\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{e}_{H,k}\right)$ $=-\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\left(\bm{I}_{H}^{h}\bm{A}\odot\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)^{*}\bm{I}_{H}^{h}\operatorname{diag}(\bm{A})\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-4\left(\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\right)^{*}\operatorname{diag}(\bm{A})\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$\leq 0\,.$
G.1 Proof for Theorem 3.3
Proof.
Solving Eqn. (17) by setting the complex gradient of the objective function to zero yields:
$\widehat{\bm{z}_{H,k}}=\bm{z}^{\prime}_{H,k}+\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}_{H}\bm{z}^{\prime}_{H,k}\right)\,,$
which, by definition, implies
$\bm{e}_{H,k}=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{z}_{H,k}^{\prime}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right)\,.$
Then, by Lemma D.1 and Lemma D.2, we have
$\bm{e}_{H,k}$ $=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{z}_{H,k}^{\prime}-\bm{I}_{h}^{H}\left(\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\bm{z}^{\prime}_{k})-\bm{Q}_{H}\bm{I}_{h}^{H}\left(\frac{\bm{W}_{\bm{z}}}{\bm{Q}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\left(\bm{z}^{\prime}_{k}-\frac{1}{\bm{Q}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\frac{\bm{W}_{\bm{z}}}{|\bm{Q}|^{2}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\frac{1}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
The proof follows from Proposition G.1 with
$\bm{A}=\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$
Appendix H Proof for Proposition 3.4
Proof.
First, if we solve for the fine-grid update at $\bm{z}_{k}^{\prime}$ in the $k$-th scanning region:
$\bm{z}_{k}^{+}=\operatornamewithlimits{argmin}_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{\prime}\right|^{2}\,,$
the resulting update is:
$\bm{e}_{k}=\bm{z}_{k}^{+}-\bm{z}_{k}^{\prime}=-\frac{1}{\bm{u}+|\bm{Q}|^{2}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\,,$
The coarse-grid update at $\bm{z}^{\prime}_{H,k}=\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\bm{z}^{\prime}_{k})$ (see Eqn. (18)) is
$\bm{e}_{H,k}=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
Hence, both fine-grid and coarse-grid updates are functions of the fine-grid gradient.
Then, with a slight abuse of notation, we define $\bm{e}_{k}(\bm{g})$ and $\bm{e}_{H,k}(\bm{g})$, which map fine-grid gradient $\bm{g}\in\mathbb{C}^{m^{2}}$ to fine-grid update in $\mathbb{C}^{m^{2}}$ and coarse-grid update in $\mathbb{C}^{(m/2)^{2}}$, respectively:
$\bm{e}_{k}(\bm{g})=-\frac{1}{\bm{u}+|\bm{Q}|^{2}}\bm{g}\quad\text{and}\quad\bm{e}_{H,k}(\bm{g})=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\bm{g}\,.$
Since $\|\bm{I}_{h}^{H}\|_{\infty}=1$ by Lemma D.4, it is straightforward to see that
$\left\{\bm{I}_{h}^{H}\bm{g}:\bm{g}\in\mathbb{C}^{m^{2}},\|\bm{g}\|_{\infty}\leq 1\right\}=\left\{\bm{g}_{H}\in\mathbb{C}^{(m/2)^{2}}:\|\bm{g}_{H}\|_{\infty}\leq 1\right\}\,.$
Hence, we observe that for all indices $r$,
$\max_{\|\bm{g}\|_{\infty}\leq 1}\left|\bm{e}_{H,k}(\bm{g})\right|_{r}$ $=\max_{\|\bm{g}_{H}\|_{\infty}\leq 1}\left|\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{g}_{H}\right|_{r}\,,$
$=\left(\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)_{r}\,,$
where the equality for all indices $r$ is achieved when $\bm{g}_{H}=\bm{J}$, a vector of ones.
Similarly, we have
$\max_{\|\bm{g}\|_{\infty}\leq 1}\left|\bm{I}_{h}^{H}\bm{e}_{k}(\bm{g})\right|_{r}=\left(\bm{I}_{h}^{H}\frac{1}{\bm{u}+|\bm{Q}|^{2}}\right)_{r}\,.$
Finally, using Lemma D.1, Lemma D.2, and Lemma D.3, we have:
$\bm{u}_{H}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\bm{u}\,,$
$\bm{u}_{H}+|\bm{Q}_{H}|^{2}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\bm{u}+|\bm{Q}_{H}|^{2}\,,$
$\bm{u}_{H}+|\bm{Q}_{H}|^{2}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\bm{u}+|\bm{Q}|^{2}\right)\,,$
$\bm{I}_{h}^{H}\left(\frac{1}{\bm{u}+|\bm{Q}|^{2}}\right)$ $=\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,,$
$\max_{\|\bm{g}_{2}\|_{\infty}\leq 1}\left|\bm{I}_{h}^{H}\bm{e}_{k}(\bm{g}_{2})\right|_{r}$ $=\max_{\|\bm{g}_{1}\|_{\infty}\leq 1}\left|\bm{e}_{H,k}(\bm{g}_{1})\right|_{r}\quad\text{for all }r\,.$